You are conducting a test of independence for the claim that there is an association between the row variable and the column variable.

X Y Z
A 30 34 21
B 40 35 21

What is the chi-square test-statistic for this data?
Answer #1

Answer:

Chi-square test-statistic = 0.777427

Solution:

Here, we have to use chi square test for independence of two categorical variables.

Null hypothesis: H0: There is no association between the row variable and the column variable.

Alternative hypothesis: Ha: There is an association between the row variable and the column variable.

We assume level of significance = α = 0.05

Test statistic formula is given as below:

Chi square = ∑[(O – E)^2/E]

Where, O is observed frequencies and E is expected frequencies.

E = row total * column total / Grand total

We are given

Number of rows = r = 2

Number of columns = c = 3

Degrees of freedom = df = (r – 1)*(c – 1) = 1*2 = 2

α = 0.05

Critical value = 5.991465

(by using Chi square table or excel)

Calculation tables for test statistic are given as below:

Observed Frequencies

Column variable

Row variable

X

Y

Z

Total

A

30

34

21

85

B

40

35

21

96

Total

70

69

42

181

Expected Frequencies

Column variable

Row variable

X

Y

Z

Total

A

32.87293

32.40331

19.72376

85

B

37.12707

36.59669

22.27624

96

Total

70

69

42

181

(O - E)

-2.87293

1.596685

1.276243

2.872928

-1.59669

-1.27624

(O - E)^2/E

0.251079

0.078677

0.08258

0.22231

0.069662

0.073118

Test Statistic = Chi square = ∑[(O – E)^2/E] = 0.777427

χ2 statistic = 0.777427

P-value = 0.677928

(By using Chi square table or excel)

P-value > α = 0.05

So, we do not reject the null hypothesis

There is not sufficient evidence to conclude that there is an association between the row variable and the column variable.